OpenStudy (anonymous):

t^(-1)(dy/dt) = 2cos^2(y) ; y(0) = pi/4

OpenStudy (anonymous):

yes, yes.... come my minions

OpenStudy (anonymous):

this is LGB's turn now :P

OpenStudy (anonymous):

oh, sorry, what?

OpenStudy (lgbasallote):

hmm \[\large t^{-1} \frac{dy}{dt} = 2\cos^2 y\] \[\large \implies \frac{dy}{tdt} = 2\cos^2y\] so you have to cross multiply...

OpenStudy (anonymous):

sooooo: 1/(2cos^2(y))dy = tdt?

OpenStudy (lgbasallote):

...well if i were you i wont move the 2...

OpenStudy (lgbasallote):

but if you wish...that works too

OpenStudy (anonymous):

i think it is more complicated than it looks like :(((

OpenStudy (lgbasallote):

actually...it's not @sami-21 lol

OpenStudy (anonymous):

no, no, it looks easier if i leave the 2

OpenStudy (lgbasallote):

okay let's leave the 2 \[\large \implies\frac{dy}{\cos^2 y} = 2tdt\] now...recall your trig identities...what is the value of \[\frac{1}{\cos^2 y}\]

OpenStudy (anonymous):

ooooone sec

OpenStudy (anonymous):

is this an inverse trig func?

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

i think it's called reciprocal trig function or something...

OpenStudy (lgbasallote):

http://www.regentsprep.org/Regents/math/algtrig/ATT1/trigsix.htm scroll down to the part it says "reciprocal functions"

OpenStudy (anonymous):

sec^2(y)?

OpenStudy (lgbasallote):

yup. \[\large \implies \sec^2 y dy = 2tdt\] now integrate both sides

OpenStudy (anonymous):

you got it :)))

OpenStudy (anonymous):

dam, so i still need to integrate...

OpenStudy (lgbasallote):

me? oh stop it youuu :""">

OpenStudy (lgbasallote):

hah cant escape integration in differential equations @IStutts :p hahaha

OpenStudy (lgbasallote):

except for higher order linear de...yeah...that has few integrations

OpenStudy (anonymous):

so tan y = t^2 + c?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

okay, now how would I continue from there?

OpenStudy (anonymous):

or at least solve for y

OpenStudy (lgbasallote):

...dont you stop there?

OpenStudy (lgbasallote):

that's rarely done in these cases...but...if you insist...

OpenStudy (lgbasallote):

take the arctan of both sides

OpenStudy (anonymous):

so y = arctan(t^2 + c)

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

and yea, my teacher is a jerk

OpenStudy (lgbasallote):

they all are =))

OpenStudy (anonymous):

so how would i solve for the initial value y(0) = pi/4?

OpenStudy (lgbasallote):

that means when t = 0 then y = pi/4

OpenStudy (anonymous):

so how did the back of the book get y = arctan ( t^2 + 1 )?

OpenStudy (lgbasallote):

because c = 1

OpenStudy (anonymous):

yea, that's how picky my teacher is, and we spent one day one separable equations, this section at least

OpenStudy (anonymous):

just by chance? or is there math i a missing?

OpenStudy (lgbasallote):

y(0) = pi/4 \[\tan y = t^2 + c\] \[\tan (\pi/4) = 0^2 + c\] \[\1 = c\]

OpenStudy (anonymous):

....oookay

OpenStudy (lgbasallote):

you didnt get it..did you...

OpenStudy (anonymous):

i never memorized the unit circle....

OpenStudy (lgbasallote):

oh lol

OpenStudy (lgbasallote):

that's gonna be your undoing now :p mwahaha

OpenStudy (anonymous):

yea, it's sad to blame my teacher, but i can only blame myself. in high school, my math teachers said it was rare they would use it, then in college....well, here i am...

OpenStudy (anonymous):

i see.... ....said the blind man

OpenStudy (lgbasallote):

well it's not too late now

OpenStudy (anonymous):

eh, i dont think my teacher will be that big of a douche. there will only be 2 of 8 questions from this chapter...

OpenStudy (anonymous):

true true

OpenStudy (lgbasallote):

maybe this can help you openstudy.com/updates/4f9be55ee4b000ae9ed11844

OpenStudy (anonymous):

anyways, onto the next one, i think i got 2.2 now

OpenStudy (lgbasallote):

http://openstudy.com/updates/4f9be55ee4b000ae9ed11844

OpenStudy (anonymous):

thank you